OpenStudy (anonymous):

Mr.harris earns $12 per hour for 40 hours of work. He earns $18 per hour for hours worked over 40. How many hours did he work if he earned $642 ?

OpenStudy (anonymous):

40 x 12 = 480 642 - 480 = 162 162/18 = 9 so he worked 40 hrs reg pay and 9 hours over time pay

OpenStudy (anonymous):

12x +18y = 642 x = 40 hours (given) 12(40) + 18y = 642 480 + 18y = 642 18y = 642-480 18y = 162 y = 162/18 = 9